(1 point) You are to roll a fair die n = 108 times, each time observing if the topside of the die shows a 6 (success) or not Heed help with c only.
Answer #1

CC) pl Between 13.). and 21.7.) PC OB < X < 0.21) = P(X-MP L2 C Xp-Up ) = P(0.13.07.2014 644 <2< 03.17 06667 6 P10.13.16 <7<0